# Top-Theta Attention: Sparsifying Transformers by Compensated Thresholding

### v1

@@ 3.1 Threshold Calibration
Figure 3 visualizes the threshold values of the 11th layer of LLaMA2-7b, which were calibrated for $k=64$. It is worth noticing that the different heads (different colors) have distinct threshold values, hence the importance of calibrating per-head thresholds. Secondly, thresholds obtained from the $A$ matrix approach a constant value as the sequence length increases, whereas thresholds obtained from the $S$ matrix tend to decrease in the longer sequence lengths. The latter is an effect of the softmax normalization that reduces the share of essential tokens out of the total row sum of 1 as the sequence length increases; therefore, keeping them requires lowering the threshold. The noisiness of the thresholds in the very last calibrated sequence length is purely due to the scarcity of the calibration samples in the dataset that corresponded to these sequence lengths (lower purple bars in Figure 3). For more examples of threshold values, refer to Appendix B.
Threshold function fitting
Our main proposal is to deploy the calibrated thresholds as complementary model parameters (one threshold parameter per every layer, head, and sequence length). However, as can be observed in Figure 3, the thresholds have a rather regular behavior as a function of the sequence length (i.e., of the attention row id). We can fit a function to it and thereby parameterize hundreds of thresholds by a few parameters.
Multi-k calibration
It is possible to calibrate multiple threshold sets $\theta_{l}(k_{1}),\theta_{l}(k_{2}),...$ targeted at different number of top elements $k_{1},k_{2},...$ to be preserved. Such calibration will allow one to dynamically decide at inference time on the desired accuracy-performance tradeoff by applying a set of thresholds of a different $k$, an attractive capability for LLM serving (Miao et al., 2023). Instead of applying  Algorithm 1 several times with the different $k$ values, we have developed a generalized calibration algorithm, which performs a single pass over the calibration samples. The generalized procedure collects up to $n$ thresholds for every attention vector of length $n$ in the calibration sample, merges these thresholds information, and eventually outputs for each (layer, head, row) a function $\theta(k)$ from $k$ to a threshold that satisfies it. Although such a calibration procedure poses large memory requirements, reducing its intensity by sampling fewer than $n$ thresholds per vector is possible. (see Appendix C).
3.2 Top-{greek}j Attention Inference
The inference procedure is depicted in Figure 1, focusing on a single transformer layer $l$, single head $h$, and single attention row $n$ of length $n$ (as in generative decoding). The attention vector is compared against the calibrated threshold value $\theta_{l,h,n}$. Attention elements that do not pass the threshold are discarded. Instead of multiplying by the entire $V$ matrix, only the selected $\tilde{k}$ row indices are loaded to a compact matrix $\tilde{V}$, which is used to compute the final product $p$. The calibrated thresholds can be stored alongside the model parameters. Since the calibration might not cover the entire range of sequence lengths $r\in\{k,k+1,\ldots\}$, a threshold of the nearest calibrated sequence length ($r$) can be used.
Applying a threshold calibrated towards a specific $k$ allows selecting $k$ attention row elements on average because the distribution of the attention values slightly varies across the processed inputs. To address the cases when thresholding results in more attention elements than desired ($\tilde{k}>k$), we experimented with capping the number of selected elements to at most k, prioritizing first and last tokens (we dubbed this technique CAPK). However, in generative tasks, CAPK caused a noticeable degradation in performance.
We propose two compensation methods that recover the accuracy degradation incurred by the Top-{greek}j by providing a better mathematical approximation of the overall self-attention operation. These methods introduce a minor additional compute but may be crucial for accuracy-sensitive tasks. Moreover, these compensations apply to other attention sparsification approaches such as the Top-k attention.
3.2.1 Softmax Denominator Compensation
Since the post-softmax sparsification showed higher accuracies in our experiments compared to pre-softmax, we are interested in approximating post-softmax-sparsified attention ($\tilde{\bm{s}}$) using pre-softmax-sparsified attention ($\tilde{\bm{a}}$). Let $\mathcal{I}\subseteq\left\{0,\dots,n-1\right\}$ denote a set of indices that we intend to keep during the sparsification, and let us establish the relation between post-softmax sparsified vector $\tilde{\bm{s}}$ and the pre-softmax sparsified vector that underwent softmax ($\mathrm{softmax}(\tilde{\bm{a}})$). Let $R$ and $E$ denote sums of exponents of the selected and discarded elements, respectively.
$\forall i\in\mathcal{I},\tilde{\bm{s}}_{i}$ $=\mathrm{softmax}(\bm{a})_{i}=\frac{e^{a_{i}-\max_{l}{a_{l}}}}{\sum_{j=0}^{n-1}e^{a_{j}-\max_{l}{a_{l}}}}$ (5)
$=\frac{e^{a_{i}-\max_{l}{a_{l}}}}{\sum_{j\in\mathcal{I}}e^{a_{j}-\max_{l}{a_{l}}}+\sum_{j\notin\mathcal{I}}e^{a_{j}-\max_{l}{a_{l}}}}$ (6)
$=\frac{e^{a_{i}-\max_{l}{a_{l}}}}{R+E}=\frac{e^{a_{i}-\max_{l}{a_{l}}}}{R}\cdot\frac{R}{R+E}$ (7)
$=\mathrm{softmax}(\tilde{\bm{a}})_{i}\cdot\frac{R}{R+E}$ (8)
In other words, to achieve a post-softmax sparsification effect, one can perform pre-softmax sparsification, estimate $\tilde{E}\approx E$, and compensate by multiplying by a factor of $R/(R+\tilde{E})$. The multiplication step can be applied after the softmax (3) or even after the SV product (4) similarly to the flash-attention (Dao et al., 2022). We consider 3 estimations:
@@ 3.2.1 Softmax Denominator Compensation
    offline-calibrated: calibrate the compensation value $\tilde{E}$ for every (layer, head, row) similarly to the method of threshold calibration.
@@
    exact: compute the exact $\tilde{E}=E$ using the sum of exponents of the non-selected row elements.
@@ 3.2.2 V-Mean-Compensation (VMC)
Applying Top-{greek}j in the following two cases will result in attention vector rows not summing up to 1: (i) post-softmax, (ii) pre-softmax followed by SDC. As a result, any attention element dropped due to thresholding will cancel the effect of its product by the corresponding row of the $V$ matrix in the final $p=sV$ product (4). In these situations, we propose compensating for the missing value by adding a mean row $\bm{\mu}$ (9) of the $V$ matrix scaled by the sum of all the discarded attention elements $\beta$ (10). Let $\bm{\tilde{\bm{s}}}\in[0,1]^{\tilde{k}}$ be a post-softmax attention vector that underwent Top-{greek}j and is eligible for VMC, and let $\bm{s}\in[0,1]^{n}$ be the non-sparsified version of it. Equations 911 show how we compute the corresponding V-mean-compensated product $\hat{p}\approx\bm{s}V$.
$\bm{\mu}$ $=\frac{1}{n}\sum_{i=0}^{n-1}V_{i}$ (9)
$\beta$ $=1-\sum_{i=0}^{\tilde{k}-1}\tilde{\bm{s}}_{i}$ (10)
@@
We justify the VMC approximation in Appendix D. The intuition behind it is to approximate each of the $n-\tilde{k}$ discarded values by an average and to multiply it by an average row of the V matrix. Such compensation can improve as the sequence length $n$ increases because the number of the averaged elements in (9) and (10) will increase accordingly. During generative decoding, $\bm{\mu}$ can be maintained as a running-mean to avoid full recalculation (9).
@@ 4 Evaluations
We use the LM Evaluation Harness (Gao et al., 2024) to evaluate the normalized accuracy metric on multiple-choice Q&A datasets, including their standard few-shot settings. In addition, we assess the Human-eval dataset using the official library (Chen et al., 2021) to measure the pass@1 metric on all 164 code generation tasks. Our source code is available [^1] https://github.com/kostyanoob/top-theta-attention.
We evaluate 3 main attention variants: (i) Baseline – full attention, without any sparsification; (ii) Top-k – keep $k$ attention elements per row; (iii) Top-{greek}j – keep $\tilde{k}\approx k$ attention elements by thresholding. We apply our compensation methods also to the Top-k baselines for a fair comparison.
@@ 4.1 Top-{greek}j Overall Performance
Figure 4: Prefill-based tasks - Tradeoff between model accuracy averaged across test samples (y-axis) and the portion of kept attention elements per attention head (x-axis). All post-softmax Top-k and Top-{greek}j employ VMC, and all pre-softmax variants employ both VMC and exact SDC. These compensations achieve little, if any, accuracy degradation while achieving up to 10× reduction in the attention elements.
This section evaluates the best configurations of Top-{greek}j, comparing them to the non-sparsified baseline and to Top-k.
@@
[…]
Overall, both Top-{greek}j and Top-k performed similarly well on both Q&A and generative tasks. For a version of Figures 4 and 5 with standard deviations of the collected metrics, the reader can refer to Appendix E. The remaining subsections closely evaluate the various aspects of our method.
(a) LLaMA3-8B-Instruct - V-rows at decoding
(b) LLaMA3-70B-Instruct V-rows at decoding
Figure 5: Generative Task Human-eval – Tradeoff between model accuracy averaged across test samples (y-axis) and the portion of required V-rows per group of heads (x-axis). The Top-{greek}j variants employ threshold calibrated on ARC-C dataset. All post-softmax Top-k and Top-{greek}j employ VMC, and all pre-softmax variants employ both VMC and exact SDC. 3× reduction of V rows is achieved.
4.2 Pre- vs Post-Softmax Thresholding
We evaluate the impact of attention matrix sparsification in its two main variants: on matrix $A$ and on the post-softmax matrix ($S$), where each of these two variants requires individual calibration. Figure 6 depicts how the different thresholdings impact the accuracy of LLaMA2-7b on the Hellaswag dataset. For comparison, the Top-k approach is also evaluated alongside Top-{greek}j. We conclude that post-softmax preserves more accuracy compared to pre-softmax thresholding. This conclusion is confirmed on extended evaluations on other datasets (see Appendix F).
Figure 6: Comparison of pre- and post-softmax sparsification
@@ 4.3 Thresholding Different Layers
We explored the impact of thresholding different layers with a different target $k$. For the LLaMA models, we have observed that targeting a slightly higher $k$ in the first layers is crucial. As a best practice, we found keeping 2 initial layers at $k=512$, whereas the rest of the layers could be sparsified more aggressively. Figure 7 shows the LLaMA2-7b model accuracy on Hellaswag, comparing Top-k and Top-{greek}j using higher $k$ in first 2 layers against using equal $k$ in all layers. All variants do not perform any compensations. The conclusion is that preserving denser attention matrices in the early layers is beneficial for the accuracy of a downstream task, which is aligned with some of the works on quantization (Tang et al., 2024; Huang et al., 2024).
Figure 7: The positive impact of high-k in first two layers
@@ 4.4 Thresholding Different Attention Rows
In Section 4.3, we have shown that Top-{greek}j attention should use individually calibrated thresholds for every transformer layer in the model due to inherently different attention element distributions across the layers, and in Section 3.1, we saw that different heads require individual thresholds. However, it is less obvious whether the different rows of the attention matrix require different thresholds within a single head. Evaluations of the LLaMA2-7b model on Hellaswag (Figure 8) show that calibrating individual thresholds per-attention row is beneficial, especially for the pre-softmax setting. For more supporting results, refer to Appendix H.
@@ 4.5 Numerical Compensations
We evaluate the proposed numerical compensation methods SDC, and VMC. Figure 9 shows that the more explicit SDC versions (exp-threshold, exact) reclaim more of the degraded accuracy on the hard task of Hellaswag. Moreover, when applying the VMC as well, the benefit is further increased.
Figure 9 focuses on the Hellaswag task, LLaMA2-7b model with Top-{greek}j attention pre-softmax, in which the non-compensated Top-{greek}j attention incurred up to 1.5% accuracy degradation, whereas the compensations have effectively closed the gap to the baseline accuracy.
In similar experiments on ARC-C and ARC-E datasets, Top-{greek}j has outperformed the baseline attention without the compensations thanks to increasing the focus of the model, or in other words, reducing the attention noise. In such cases, the compensations do not contribute much and may only bring back the noise, thereby degrading the accuracy down to the baseline accuracy levels.
Figure 9: Positive impact of SDC and VMC compensations
@@ 4.6 Impact of Grouped Query Attention (GQA)
In GQA, multiple attention heads share the same V matrix. In LLaMA-3-8B models, the GQA group size is $g=4$; therefore, when Top-k attention is applied, in the worst case of total disagreement between the heads, there could be $4\cdot k$ rows of V necessary. However, as seen in Figure 10, both Top-k and Top-{greek}j set for $k=128$ remain at approximately 250-300 selected V-rows per group. Similar results were observed in numerous layers, indicating that different heads in the group do have a certain agreement. To observe the selected tokens of the heads in the group, refer to Appendix I.
Figure 10: GQA impact - Number of required V-rows for every generated token, LLaMA-3-8B-Instruct (layer 2, first GQA group), Human-eval task #25.
In our experiments, each head in the group only used its selected attention elements and V-rows. However, we envision 2 potentially better methods to deal with GQA from a sparsification perspective: 1) discarding V-rows that were requested by the Top-{greek}j masks in very few heads in the group. 2) Augment the per-head $\mathcal{I}$ with the group’s union - to reclaim more accuracy using V rows that are loaded anyway for the group.
@@ 4.7 Distribution Shifts
To examine how resilient the calibrated threshold is, we evaluate on Human-eval the two following Top-{greek}j variants: 1) calibrated on Q&A dataset of ARC-C, and 2) calibrated on the first 10% of Human-eval tasks. As seen in Figure 11, the Top-{greek}j post-softmax is even more accurate when using thresholds calibrated on a different dataset. For pre-softmax, there is a benefit of calibrating on the same dataset. Overall, the thresholds show resilience towards distribution shift, which suggests that they are strongly associated with the model rather than with the data. This allows calibrating thresholds once per model.
Figure 11: Distribution shift LLaMA3-8b-Instruct evaluated on Human-eval. cal.arcc variants were calibrated on ARC-C

### v2

@@ 3.2.1 Softmax Denominator Compensation (SDC)
$\forall i\in\mathcal{I},\tilde{\bm{s}}_{i}$ $=\mathrm{softmax}(\bm{a})_{i}=\frac{e^{a_{i}-\max_{l}{a_{l}}}}{\sum_{j=0}^{n-1}e^{a_{j}-\max_{l}{a_{l}}}}=\frac{e^{a_{i}-\max_{l}{a_{l}}}}{\sum_{j\in\mathcal{I}}e^{a_{j}-\max_{l}{a_{l}}}+\sum_{j\notin\mathcal{I}}e^{a_{j}-\max_{l}{a_{l}}}}$
$=\frac{e^{a_{i}-\max_{l}{a_{l}}}}{R+E}=\frac{e^{a_{i}-\max_{l}{a_{l}}}}{R}\cdot\frac{R}{R+E}=\mathrm{softmax}(\tilde{\bm{a}})_{i}\cdot\frac{R}{R+E}$ (4)
[…]
@@
    offline-calibrated: calibrate a static compensation value $\tilde{E}$ for every (layer, head, row) similarly to the method of threshold calibration.
@@
    exact: compute $\tilde{E}=E$ by summing up exponents of the non-selected row elements.
@@ 3.2.2 V-Mean-Compensation (VMC)
Applying Top-{greek}j in the following two cases will result in attention vector not summing up to 1: (i) post-softmax, (ii) pre-softmax followed by SDC. As a result, the value given by the product of non-selected attention elements and their corresponding $\bm{V}$-rows will be missing in the final product (3). We denote these cases as “eligible for VMC” and we compensate for the missing value in (3) by adding a mean row $\bm{\mu}$ of the $\bm{V}$ matrix scaled by the sum of all the discarded attention elements $\beta$. Equation 7 lists the computation of the corresponding $\bm{V}$-mean-compensated product $\hat{\bm{p}}\approx\bm{s}\bm{V}$, where $\bm{\tilde{\bm{s}}}\in[0,1]^{\tilde{k}}$ is a post-softmax attention vector that underwent Top-{greek}j (containing the selected $\tilde{k}$ out of $n$ attention elements) and is eligible for VMC, and $\bm{\tilde{V}}$ contains selected $\tilde{k}$ rows of $\bm{V}$.
$\bm{\mu}$ $=\frac{1}{n}\sum_{i=0}^{n-1}\bm{V}_{i}$ (5)
$\beta$ $=1-\sum_{i=0}^{\tilde{k}-1}\tilde{\tilde{\bm{s}}}_{i}$ (6)
@@
We formally justify the VMC approximation in Appendix˜E. The intuition behind it is to approximate each of the $n-\tilde{k}$ discarded attention values by an average and to multiply it by an average row of the $\bm{V}$ matrix. Such compensation can improve as the sequence length $n$ increases because the number of averaged elements in $\bm{\mu}$ and $\beta$ will increase accordingly. During generative decoding, $\bm{\mu}$ can be maintained as a running mean to avoid full recalculation.
@@ 4 Evaluations
We use the LM Evaluation Harness Gao et al. (2024) to evaluate the normalized accuracy metric on multiple-choice Q&A datasets, including their standard few-shot settings. In addition, we assess the Human-eval dataset using the official evaluation harness Chen et al. (2021) to measure the pass@1 metric on all 164 code generation tasks, and two long sequence summarization tasks of LongBench (qmsum and gov_report) using their official evaluation setup Bai et al. (2024). Our hardware setup requires GPUs with a total of 192GB (to run 70B models), and a CPU with 300GB of physical memory. Reproducing the 7B and 8B model results on smaller datasets (i.e., not LongBench) can be accomplished using a single 24GB VRAM GPU. For detailed software versions and reproducibility scripts, we make our source code available [^1] https://github.com/kostyanoob/top-theta-attention.
We evaluate 3 main attention variants: (i) Baseline – full attention, without any sparsification; (ii) Top-k – keep $k$ attention elements per row; (iii) Top-{greek}j – keep $\tilde{k}\approx k$ attention elements by thresholding. We also apply our compensation methods to the Top-k baselines for a fair comparison.
@@ 4.1 Top-{greek}j Overall Performance
This section compares the best configurations of Top-{greek}j against the non-sparsified baseline and Top-k.
@@
Figure 4: Prefill-based tasks - Tradeoff between model accuracy (y-axis) and the portion of kept attention elements per attention head (x-axis). All post-softmax Top-k and Top-{greek}j employ VMC, and all pre-softmax variants employ both VMC and exact SDC. These compensations achieve little, if any, accuracy degradation while achieving up to 10× reduction in the attention elements.
In Figure˜4, LLaMA2 Touvron et al. (2023) and LLaMA3 Grattafiori et al. (2024) models are evaluated on Q&A tasks, where the $k$ parameter of Top-k and of the Top-{greek}j is swept from 32 to 512. The first two layers are kept at $k=512$. The calibration set used for Top-{greek}j in each dataset is $10\%$ of the training or validation sets (different from the test set). The $x$-axis shows the fraction of the attention elements that are involved in computation, normalized to an average number of attention elements in the entire model in a given forward pass. The $y$-axis shows the normalized accuracy. The main observation is that in all models, both Top-k and Top-{greek}j have increased the accuracy (by $0.2\%-1\%$) compared to the baseline while pruning away a significant portion of attention elements ($2\times$$5\times$ fewer elements were active). Secondly, post-softmax sparsification performs consistently better in both Top-k and Top-{greek}j, compared to the pre-softmax sparsification.
(a) LLaMA3-8B-Instruct - HumanEval
(b) LLaMA3-70B-Instruct - HumanEval
(c) LLaMA3.1-8B-Instruct LongBench-qmsum
(d) LLaMA3.1-8B-Instruct LongBench-gov_report
Figure 5: Generative Tasks – Tradeoff between model accuracy (y-axis) and the portion of required $\bm{V}$-rows per group of heads (x-axis). The Top-{greek}j variants employ a threshold calibrated on ARC-C dataset. All post-softmax Top-k and Top-{greek}j employ VMC, and all pre-softmax variants employ VMC and exact SDC. 3× and 10× reduction of $\bm{V}$ rows is achieved on Human-eval and LongBench, respectively.
In Figure˜5, we focus on the generative tasks, where the main bottleneck is the reading of the KV cache. To demonstrate domain adaptation capabilities of the calibrated Top-{greek}j thresholds, they were first calibrated on a different dataset (ARC-C) and then loaded for Human-eval and LongBench evaluation. We examine LLaMA-3-Instruct and LLaMA-3.1-Instruct models, and we use 0 temperature for generation.
On Human-eval task ( Figures˜5(a) and 5(b)) we show the tradeoff between the output quality (pass@1) and the number of required $\bm{V}$-rows. The $k$ parameter was swept between 32 and 512 with the first two layers set at 512. The post-softmax Top-{greek}j outperforms other approaches in both 8B and 70B models, preserving pass@1 within $1\%$ of the baseline while reducing the required $\bm{V}$-rows by $3\times$. Impressively, in the 70B model, Top-{greek}j offered a 5× reduction at the expense of $1\%$ of the pass@1.
On LongBench tasks (Figures˜5(c) and 5(d)), where the average prompt lengths reached around 15k tokens, the $k$ parameter was swept in between 128 and 768 with the first two layers set at 768, and Top-{greek}j showed even higher reductions of the required $\bm{V}$-rows - up to 10×  despite the GQA. Such a reduction was achieved using thresholds calibrated for as few as $k=128$ elements. The Rouge-L score (higher is better) stayed within $1\%$ of the baseline, often improved by $0.5\%$.
Overall, both Top-{greek}j and Top-k performed similarly well on both Q&A and generative tasks. For a version of Figures˜4 and 5 with standard deviations of the collected metrics, the reader can refer to Appendix˜F. The rest of this section shows an ablation study of our method.
4.2 Pre- vs Post-Softmax Thresholding
@@
We evaluate the impact of attention matrix sparsification in its two main variants: on matrix $\bm{A}$ and on the post-softmax matrix ($\bm{S}$), where each of these two variants requires individual calibration. Figure˜6 depicts how the different thresholdings impact the accuracy of LLaMA2-7b on the Hellaswag dataset. For comparison, the Top-k approach is also evaluated alongside Top-{greek}j. We conclude that post-softmax preserves more accuracy compared to pre-softmax thresholding, and we provide extended results on additional datasets (see Appendix˜G).
@@ 4.3 Thresholding Different Layers
Figure 7: High-k in first two layers
We explored the impact of thresholding different layers with a different target $k$. For the LLaMA models, we have observed that targeting a slightly higher $k$ in the first layers is crucial. As a best practice, we found keeping 2 initial layers at $k=512$, whereas the rest of the layers could be sparsified more aggressively. Figure˜7 shows the LLaMA2-7b model accuracy on Hellaswag, comparing Top-k and Top-{greek}j using higher $k$ in first 2 layers against using equal $k$ in all layers. All variants do not perform any compensations. The conclusion is that preserving denser attention matrices in the early layers is beneficial for the accuracy of a downstream task, which is aligned with some of the works on quantization Tang et al. (2024); Huang et al. (2024).
@@ 4.4 Thresholding Different Attention Rows
Figure 8: Calibrating per-attention-row thresholds vs a unified threshold for all attention rows
In Section˜4.3, we have shown that Top-{greek}j attention should use individually calibrated thresholds for every transformer layer in the model due to inherently different attention element distributions across the layers, and in Section˜3.1, we saw that different heads require individual thresholds. However, it is less obvious whether the different rows of the attention matrix require different thresholds within a single head. Evaluations of the LLaMA2-7b model on Hellaswag (Figure˜8) show that calibrating individual thresholds per-attention row is beneficial, especially for the pre-softmax setting. For more supporting results, refer to Appendix˜I.
@@ 4.5 Numerical Compensations
Figure 9: SDC and VMC compensations impact the accuracy positively.
We evaluate the proposed numerical compensation methods SDC and VMC, finding that more explicit SDC variants (exp-threshold, exact) substantially recover degraded accuracy on the challenging Hellaswag task with LLaMA2-7b and Top-{greek}j attention, as shown in Figure˜9; combining SDC with VMC further improves results, effectively closing the accuracy gap between the baseline and the non-compensated Top-{greek}j attention. However, on the ARC-C and ARC-E datasets, Top-{greek}j attention alone already outperforms the baseline by reducing attention noise, so applying compensations in these cases offers little benefit and may even reduce accuracy back to baseline levels by reintroducing noise. Overall, SDC and VMC compensations almost entirely recover the accuracy.
@@ 4.6 Impact of Grouped Query Attention (GQA)
Figure 10: GQA impact - Number of required $\bm{V}$-rows for every generated token, LLaMA-3-8B-Instruct (layer 2, first GQA group), Human-eval task #25.
In GQA, multiple attention heads share the same $\bm{V}$ matrix; for LLaMA-3-8B models with a group size of $g=4$, this means up to $4k$ $\bm{V}$ rows could be needed if heads select completely different tokens. However, as shown in Figure˜10, both Top-k and Top-{greek}j with $k=128$ typically select only 250–300 $\bm{V}$ rows per group, indicating substantial agreement among heads - a pattern observed across many layers and further illustrated in Appendix˜J.
In our experiments, each head in the group used only its own selected attention elements and $\bm{V}$-rows, but two alternative approaches could improve GQA sparsification: (1) discarding $\bm{V}$-rows requested by only a few heads, and (2) augmenting each head’s selections with the group’s union to leverage $\bm{V}$-rows already loaded for the group and potentially improve accuracy.
@@ 4.7 Distribution Shifts
To examine how domain-sensitive the calibrated threshold is, we evaluate on Human-eval the two following Top-{greek}j variants: 1) calibrated on Q&A dataset of ARC-C, and 2) calibrated on the first 10% of Human-eval tasks. As seen in Figure˜11(a), the Top-{greek}j post-softmax is even more accurate when using thresholds calibrated on a different dataset. For pre-softmax, there is a benefit of calibrating on the same dataset. We also evaluate using MedMCQA, where we compare the Top-{greek}j calibrated on ARC-C with Top-k on Figure˜11(b) - showing that both sparsification methods perform equally well. Overall, the thresholds show resilience towards distribution shift, suggesting that they are strongly associated with the model rather than with the data. This allows calibrating thresholds once per model.
(a) LLaMA3-8B-Instruct - Human-eval
(b) LLaMA3-8B MedMCQA
Figure 11: Distribution shift - Top-{greek}j calibrated on different task (labeled with cal.arcc) shows comparable accuracy and attention reduction compared to Top-{greek}j calibrated on the same task 11(a), and compared to Top-k 11(b).